# Reflect3r: Single-View 3D Stereo Reconstruction Aided by Mirror Reflections

### v1

@@ Abstract
Mirror reflections are common in everyday environments and can provide stereo information within a single capture, as the real and reflected virtual views are visible simultaneously. We exploit this property by treating the reflection as an auxiliary view and designing a transformation that constructs a physically valid virtual camera, allowing direct pixel-domain generation of the virtual view while adhering to the real-world imaging process. This enables a multi-view stereo setup from a single image, simplifying the imaging process, making it compatible with powerful feed-forward reconstruction models for generalizable and robust 3D reconstruction. To further exploit the geometric symmetry introduced by mirrors, we propose a symmetric-aware loss to refine pose estimation. Our framework also naturally extends to dynamic scenes, where each frame contains a mirror reflection, enabling efficient per-frame geometry recovery. For quantitative evaluation, we provide a fully customizable synthetic dataset of 16 Blender scenes, each with ground-truth point clouds and camera poses. Extensive experiments on real-world data and synthetic data are conducted to illustrate the effectiveness of our method.
@@ 2.2 Single-Image-to-3D Reconstruction
Generative. The first line recovers 3D geometry through generation by hallucinating multi-view images of a 3D model and reconstructing the scene from the synthesized dense views. Early approaches, such as RealFusion [22] and Make-It-3D [30], adopt a per-scene optimization strategy, distilling prior knowledge from 2D generative models into 3D representations via the Score Distillation Sampling (SDS) loss [27]. While initially designed for object-level reconstruction, this paradigm has been extended to scene-level synthesis [37, 36]. More recent methods [20, 45, 16, 10] bypass per-scene optimization by directly training diffusion models that, conditioned on an input view and target camera parameters, predict novel views which are then used for 3D reconstruction.

### v2

@@ Abstract
Mirror reflections are common in everyday environments and can provide stereo information within a single capture, as the real and reflected virtual views are visible simultaneously. We exploit this property by treating the reflection as an auxiliary view and designing a transformation that constructs a physically valid virtual camera, allowing direct pixel-domain generation of the virtual view while adhering to the real-world imaging process. This enables a multi-view stereo setup from a single image, simplifying the imaging process, making it compatible with powerful feed-forward reconstruction models for generalizable and robust 3D reconstruction. To further exploit the geometric symmetry introduced by mirrors, we propose a symmetric-aware loss to refine pose estimation. Our framework also naturally extends to dynamic scenes, where each frame contains a mirror reflection, enabling efficient per-frame geometry recovery. For quantitative evaluation, we provide a fully customizable synthetic dataset of 16 Blender scenes, each with ground-truth point clouds and camera poses. Extensive experiments on real-world data and synthetic data are conducted to illustrate the effectiveness of our method. Code and Blender scenes are available in: https://github.com/jingwu2121/reflect3r.git.
@@ 2.2 Single-Image-to-3D Reconstruction
Generative. The first line recovers 3D geometry through generation by hallucinating multi-view images of a 3D model and reconstructing the scene from the synthesized dense views. Early approaches, such as RealFusion [22] and Make-It-3D [30], adopt a per-scene optimization strategy, distilling prior knowledge from 2D generative models into 3D representations via the Score Distillation Sampling (SDS) loss [27]. While initially designed for object-level reconstruction, this paradigm has been extended to scene-level synthesis in works like Reconfusion [36]. More recent methods [20, 44, 16, 10] bypass per-scene optimization by directly training diffusion models that, conditioned on an input view and target camera parameters, predict novel views which are then used for 3D reconstruction.
@@ 6 Conclusion
Supplementary Material
7 Implementation Details
We explain more details about the mirror plane recovery in Sec. 5.1. Since DUSt3R lacks direct correspondences in the mirror region, its predicted mirror plane is often inaccurate in position, though the normal direction is generally reliable. Our objective is to recover both the plane’s normal and its position. Our goal is to obtain the normal and the position of the plane. The normal is estimated using the method described in Eq. (11). For plane positioning, we first extract edge points in the main view via image-space edge detection and back-project them into 3D. RANSAC is then applied to remove outliers. To improve robustness, we retain the top 10% of edge points ranked by DUSt3R ’s confidence scores and randomly select one as a reliable anchor point to finalize the plane position.
8 More Details About the Synthetic Data
We provide the thumbnail of all 16 scenes included in our dataset in Fig. 10.
Figure 9 shows an example Blender scene in our dataset. Each scene in the dataset can be fully customized, including object shapes, room layouts, furniture placement, material properties, lighting conditions, etc. We additionally insert and adjust mirror surfaces with controllable positions and reflectance properties to simulate realistic reflective setups. This design not only ensures diverse and detailed scenes for training and evaluation but also provides a flexible foundation for extending the dataset in future work.
Figure 9: Details of the synthetic dataset’s Blender modeling. Top left: wireframe view; top right: solid view; bottom left: material view; bottom right: rendered view. The scene is fully customizable, allowing adjustments to the mirror’s position and properties, as well as the room setup and lighting, enabling easy extension of the dataset.
Figure 10: Thumbnails of the 16 fully editable Blender scenes included in our dataset.
9 More Details About the Experiments
9.1 Details of Evaluation Metrics
We cover the math of the evaluation metrics in this section.
Given the ground-truth point cloud $\mathcal{P}=\{\mathbf{p}_{i}\}_{i=1}^{N_{P}}$ and the reconstructed point cloud $\mathcal{Q}=\{\mathbf{q}_{i}\}_{i=1}^{N_{Q}}$, Chamfer Distance is computed as
[…]
In our paper, we report the completeness (Comp.) and the accuracy (Accu.) in percentage with a threshold of 1 cm of the point cloud distance. Specifically, let the indicator function be $\mathbf{l}[\cdot]$,
[…]
F1 $=\frac{2\cdot\text{Comp.}\cdot\text{Accu.}}{\text{Comp.}+\text{Accu.}}.$ (20)
Figure 11: Boxplots of completeness, accuracy, and F1 score on the synthetic dataset.
9.2 More Statistics And Analysis
To further assess robustness, we visualise the score distributions in Fig. 11 using boxplots for completeness, accuracy, and F1 score on the synthetic dataset. Reflect3r achieves the highest median scores across all metrics while also exhibiting the smallest interquartile range and whisker span, indicating low variance and strong stability across scenes. MoGe attains similar median scores to Reflect3r, suggesting that its predicted point clouds are close to the ground-truth, but its results fluctuate significantly, reflecting high variance. Its completeness is notably lower, as MoGe does not leverage the stereo information provided by mirrors and therefore fails to recover occluded regions, leaving parts of the scene uncovered. DUSt3R and VGGT, unable to identify mirrors, generate false geometry in reflective regions and consequently show much higher variability. MASt3R, which cannot handle single-view reconstruction, consistently predicts flat geometry and therefore yields uniformly low scores.